# A note on solving basic equations over the semiring of functional digraphs

### v1

@@ 4 Conclusion
This paper proves that DEEP  and some of its variants are in ¶. For some other variants we were able to prove only polynomially testable necessary conditions. These results will greatly improve the performances of the software pipeline for solving general polynomial equations over functional graphs proposed in Dennunzio et al. (2023).
Several improvements are possible. The first one is to reproduce similar results for general functional graphs. A second one would investigate effective algorithms to enumerate solutions of basic equations in an effective way i.e. in such a way that producing two successive solutions is polynomial in the size of the first solution.

### v3

@@ 4 Conclusion
This paper proves that DEEP  and some of its variants are in ¶. For some other variants, we were able to prove only polynomially testable necessary conditions. These results will greatly improve the performances of the software pipeline for solving general polynomial equations over functional graphs proposed in [4] when one is interested in the existence of solutions and not to their enumeration.
Several improvements to our results are possible. The first one is, of course, to reproduce similar results for general functional graphs, taking into accounts new recent results from [5] and [8].
@@
Acknowledgements
This work was partially supported by:
  • 1.
    the HORIZON-MSCA-2022-SE-01 project 101131549 “Application-driven Challenges for Automata Networks and Complex Systems (ACANCOS)”;
  • 2.
    the PRIN 2022 PNRR project “Cellular Automata Synthesis for Cryptography Applications (CASCA)” (P2022 MPFRT) funded by the European Union – Next Generation EU;
  • 3.
    ANR “Ordinal Time Computations” (OTC), ANR-24-CE48-0335-01.